OpenStudy (anonymous):

Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Provide the solution to the system. 5x - y = -2 10x - 3y = -7

OpenStudy (anonymous):

choose an equation to set y equal to ex: 5x-y=-2 can also be in the form of y= 5x+2 plug in y=5x+2 into 10x-3y+-7 10x-3(5x+2)=-7 -5x-6=-7 -5x=-1 x=1/5 now plug the sol x into x in one of the equation to get y 10(1/5)-3y=-7 2-3y=-7 -3y=-9 y=3 The solution is (1/5,3)